OpenStudy (anonymous):

lim as x=>3- ((sqrtx^2+16)-5)/(x^2-9)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3-}\] \[{\sqrt{x^2+16}-5}/{\sqrt{x^2-9}}\]

OpenStudy (nikvist):

\[x\rightarrow 3-\,\Rightarrow\sqrt{x^2-9}\,\notin\,\mathbb{R}\]

OpenStudy (anonymous):

right.

OpenStudy (nikvist):

then, limit does not exist

OpenStudy (anonymous):

there's no way to factor out the discontinuity?